# Lagrangian neural networks for nonholonomic mechanics

### v1

@@ 7 Conclusions
Regarding the loss graphs for training and testing for both models in the examples, we observe that although the initial values are nearly identical, the LNN-nh model shows a significantly steeper decrease in the loss function during training and testing. By the end of training, the LNN-nh model’s loss is two orders of magnitude lower than that of the LNN model. This difference in loss is evident in the greater deviation of the LNN model’s predicted accelerations from the actual values, as shown in the corresponding graphs for each example.
@@
The examples show that our model achieves significantly lower loss in both the training and test sets than the LNN model. It also consistently demonstrates effective energy stability and conservation, maintaining nearly constant energy levels along different trajectories. The implementations indicate that constraints from the LNN-nh model are more robust and stable over time than those from the LNN model, which are more susceptible to changes and rapidly drift away from the initial zero value of the constraint.
@@
To summarize, across all experiments, the networks that incorporate the nonholonomic treatment of constraints into the loss function generally outperform the Lagrangian neural networks that do not consider the non-holonomic constraints. The results highlight the effectiveness of incorporating nonholonomic constraints in improving neural network performance for systems with such kind of restrictions.

### v2

@@ 7 Conclusions
Regarding the loss graphs for training and testing for both models in the examples, we observe that although the initial values are nearly identical, the LNN-nh model shows a significantly steeper decrease in the loss function during training and testing in linear cases, and almost the same performance in the nonlinear case. By the end of training, the LNN-nh model’s loss is two orders of magnitude lower than that of the LNN model in examples with linear constraints. This difference in loss is evident in the greater deviation of the LNN model’s predicted accelerations from the actual values, as shown in the corresponding graphs for each example.
@@
The examples show that our model achieves equal or lower loss in both the training and testing sets than the LNN model. It also consistently demonstrates effective energy stability and conservation, maintaining nearly constant energy levels along different trajectories. The implementations indicate that constraints from the LNN-nh model are more robust and stable over time than those from the LNN model, which are more susceptible to changes and rapidly drift away from the initial zero value of the constraint.
@@
To summarize, across all experiments, the networks that incorporate the nonholonomic treatment of constraints into the loss function generally outperform the Lagrangian neural networks that do not consider the non-holonomic constraints. The results highlight the effectiveness of incorporating nonholonomic constraints in improving Lagrangian neural network performance for systems with such kind of restrictions.